# Mean-field Potts and random-cluster dynamics from high-entropy initializations

### v1

@@ 4.1 Preliminaries for the Potts Glauber dynamics
    $\exists\delta\geq 0$: $\mathbf{E}_{x_{0}}[X_{t+1}-X_{t}\mid\mathcal{F}_{t}]\leq-\delta$ on $\{X_{t}\geq 0\}$ for all $t\geq 0$.
@@ 4.2 Mixing away from the saddle when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$
We first establish that as soon as the proportion vector has reached an $\omega(n^{-1/2})$ distance from the unstable fixed point of $(m_{*},\frac{1-m_{*}}{q-1},...,\frac{1-m_{*}}{q-1})$ in the first coordinate, with probability going to $1$ as $n\to\infty$, the Glauber dynamics rapidly quasi-equilibrates to the corresponding phase. We need to handle the regimes $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$ and $\beta>\beta_{\mathsf{s}}$ separately, with the latter introducing additional complications; this subsection is focused on the former. Define the $q+1$ stable Potts phases when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$ as
$\Omega^{\mathsf{dis}}$ =$=\{\sigma:\|S(\sigma)-(\tfrac{1}{q},...,\tfrac{1}{q})\|_{1}\leq o(n)\}\,,$
$\Omega^{{\mathsf{ord}},i}$ =$=\{\sigma:\|S(\sigma)-(\tfrac{1-m_{\mathsf{r}}}{q-1},...,\tfrac{1-m_{\mathsf{r}}}{q-1},m_{\mathsf{r}},\tfrac{1-m_{\mathsf{r}}}{q-1},...,\tfrac{1-m_{\mathsf{r}}}{q-1})\|_{1}\leq o(n)\}\,,\qquad\text{and}$
@@
The following allows us to treat the $q-1$ non-dominant coordinates as constant $\frac{1-S_{t,1}}{q-1}$ and approximate the analysis by a 1-dimensional process for the first coordinate.
@@
$\mathbb{P}(\tau_{i,j}^{+}\leq T\wedge\tau_{\varepsilon})\leq 2\exp\Big{(}-\frac{(\frac{A_{2}\log n}{\sqrt{n}}-2/n)^{2}}{8T\cdot(2/n)^{2}}\Big{)}\leq 2\exp\Bigg{(}-\frac{\big{(}\frac{A_{2}\log n}{2\sqrt{n}}\big{)}^{2}}{\frac{2A_{1}n\log n}{n^{2}}}\Bigg{)}=2\exp\Bigg{(}-\frac{A_{2}^{2}\log n}{8A_{1}}\Bigg{)}\,.$
@@
$\left(\frac{d}{dx_{i}}g_{\beta,i}\Big{|}_{x=\hat{S}_{t}}-\frac{d}{dx_{i}}g_{\beta,j}\Big{|}_{x=\hat{S}_{t}}\right)=\frac{\beta e^{\beta\cdot\hat{S}_{t,i}}}{\sum_{k=1}^{q}e^{\beta\cdot\hat{S}_{t,k}}}$ $<\frac{qe^{\beta\cdot\hat{S}_{t,i}}}{\sum_{k=1}^{q}e^{\beta\cdot\hat{S}_{t,k}}}$
@@
If $W_{t}\geq 0$, then the first term is at most $-\frac{\eta Kn^{-1/2}\log n}{n}$ and the second term is lower order, yielding (60). ∎
@@
$=\frac{1}{n}\cdot[D_{\beta}(m_{*})+(S_{t+\tau_{i},1}-m_{*})\cdot c_{*}+O(|S_{t+\tau_{i},1}-m_{*}|^{2})]+O(n^{-2})$ (67)
@@
where the steps in the last line of (68) hold since $-s(1+\frac{c_{*}}{16})^{2}<S_{t+\tau_{i},1}-m_{*}<0$, and $s$ is sufficiently small in terms of $c_{*}$ that the second order term is at most $c_{*}(S_{t+\tau_{i},1}-m_{*})/2$, say.
@@
$\big{\|}S(\sigma_{0})-\big{(}m_{\mathsf{r}},\tfrac{1-m_{\mathsf{r}}}{q-1},...,\tfrac{1-m_{\mathsf{r}}}{q-1}\Big{)}\big{\|}_{1}\leq\frac{r}{\sqrt{n}}\,,\qquad\text{and}\qquad\big{\|}S(\sigma_{0}^{\prime})-\big{(}m_{\mathsf{r}},\tfrac{1-m_{\mathsf{r}}}{q-1},...,\tfrac{1-m_{\mathsf{r}}}{q-1}\Big{)}\big{\|}_{1}\leq\frac{r}{\sqrt{n}}\,.$
@@
$\mathbb{P}\left(\tau_{i+1}\leq\tau_{i}+n,\tilde{\tau}_{i+1}>\min\{\tau_{k+1},\tau_{i}+n\},\tau^{+}_{i+1}>\min\{\tau_{k+1},\tau_{i}+n\}\mid S_{\tau_{i}}\right)\geq 1-3r_{i}-O(n^{-2}).$ (71)
@@ 4.3 Getting away from the saddle point when $\beta=\beta_{c}$
$\mathbb{E}[Z_{t+1}-Z_{t}\mid\mathcal{F}_{t}]=n\mathbb{E}[S_{t,1}-S_{t+1,1}\mid\mathcal{F}_{t}]=-d_{\beta}(S_{t})+O(n^{-1})\geq-D_{\beta}(S_{t,1})+O(n^{-1})\geq-C_{1}\gamma n^{-1/2},$
@@
Consider the time and space rescaled process $\bar{S}_{t}=\sqrt{n}(S_{tn}-(m_{*},\frac{1-m_{*}}{q-1},...,\frac{1-m_{*}}{q-1}))$. By standard results regarding limits of discrete stochastic dynamics as stochastic differential equations (see, e.g., [37]), $\bar{S}_{t}$ converges to the solution of a stochastic differential equation (SDE) $Z_{t}$ on $\mathcal{S}$ with drift and volatility functions with bounded Lipschitz constants. On a compact space, like $\mathcal{S}$, such convergence results only require $O(1)$-Lipschitz bounds on the drift function (which hold in our setting for $g_{\beta}-s$) and moment estimates on the step-wise increments (for which ours are uniformly bounded by ±$\pm\frac{1}{\sqrt{n}}$).
@@
If for each initial $\hat{c}_{*}$ we define $f(\hat{c}_{*})$ to be the probability that $\bar{S}_{t,1}$ hits $-\gamma$ before $\gamma$ and we define $f^{Z}(\hat{c}_{*})$ to be the same probability for $Z_{t}^{1}$< then by the convergence described above,
@@
for all $\hat{c}_{*}$. Furthermore, $f^{Z}$ is easily checked to be $n$-independent, continuous, monotone, and going to 00 as $\hat{c}_{*}\to\gamma$ and $1$ as $\hat{c}_{*}\to-\gamma$. Therefore, there is a unique $\hat{c}_{*}$ where $f^{Z}$ is the desired $\xi$. Finally, to see that the choice of $\hat{c}_{*}$ is $\gamma$-independent up to $o_{\gamma}(1)$, note that if $\tau_{\gamma}^{-}<\tau_{\gamma}^{+}$, then with probability $1-o_{\gamma}(1)$, also $\tau_{2\gamma}^{-}<\tau_{2\gamma}^{+}$ (by the argument in Lemma 4.6). ∎
@@ 4.4.1 Fixed point analysis of the $q$-dimensional drift function
Now, by (59), $D_{\beta}^{\prime}(x)=0$ has at most two zeros. Aside from $x^{+}$, let $x^{++}$ denote the other zero (if it exists). If $x^{++}$ does not exist or $x^{++}\leq m_{\mathsf{r}}$, then $D_{\beta}^{\prime}(x)<0$ for all $x>m_{\mathsf{r}}$ so $m_{\mathsf{r}}$ is the unique zero in $(1/q,1)$. Now assume $x^{++}>m_{\mathsf{r}}$. By continuity of $D_{\beta}^{\prime}$, $D_{\beta}^{\prime}(x)<0$ for all $x\in(m_{\mathsf{r}},x^{++})$, so by integration $D_{\beta}(x)<0$ for $x\in(m_{\mathsf{r}},x^{++}]$. Moreover, note that $D_{\beta}$ is a monotone function in $(x^{++},1]$ since $D_{\beta}^{\prime}$ does not longer change sign in this interval. Hence, $D_{\beta}$ has no zero on $[x^{++},1]$ and $m_{\mathsf{r}}$ is the unique root in $(1/q,1]$. ∎
@@
We first reason that all solutions are of the form of perturbations of $(a,...,a,\frac{1-ka}{q-k},...,\frac{1-ka}{q-k})$. In order to see this, we suppose by way of contradiction there exists a solution $\bar{s}$ having three distinct values $a,b,c$ appearing in its proportions vector (wlog as $\bar{s}_{1}=a,\bar{s}_{2}=b$ and $\bar{s}_{3}=c$; let $\bar{Z}=\sum_{j}e^{\beta\bar{s}}$). Then, $e^{\beta a}=a\bar{Z}$ and similarly for $b$ and $c$, and $a\neq b\neq c$. But this is impossible because for any $C,\beta>0$, the equation $e^{\beta x}=Cx$ has at most two solutions on $[0,1]$. When $k=1$, the only possible such solution is the one which has $D_{\beta}=0$ (because the other coordinates are equal, which is where $D_{\beta}=d_{\beta}$), and we can apply Lemma 4.17 to read off that in that case $a=m_{\mathsf{r}}$.
@@
as that would say that such a fixed point is unstable with a drift towards the first coordinate increasing if we perturb in the $(1,-1,0,...,0)$ direction. To see this, differentiating $g_{\beta,1}$ and plugging in $g_{\beta,1}(\bar{s})=\bar{s}$ and $\bar{s}_{1}=\bar{s}_{2}=\bar{a}$,
@@ 4.4.2 Getting to the stable fixed point once one coordinate dominates
Suppose $S_{0}$ is such that $S_{0,1}\geq S_{0,i}+\delta$ for all $i\geq 2$ for some $\delta>0$. Then, for every $\rho>0$, there exists a time $T=O(n)$ such that $\|S_{T}-(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})\|_{1}\leq\rho$ with probability $1-o(1)$.
@@
therefore it has no closed orbits, and by the bounded Lipschitz constant of $d_{\beta}$, converges exponentially fast to the fixed point in whose basin it is initialized. Moreover, by Lemma 4.18, if initialized with $\mathbf{S}_{0,1}\geq\mathbf{S}_{0,i}+\delta$ for all $i=2,...,q$ for some $\delta>0$, then it is in the basin of attraction of the stable fixed point $(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})$. In particular, for every $\delta>0$, and $\rho>0$, there is a $T=O(1)$ such that under the assumptions on $\mathbf{S}_{0}$ of the lemma, the system of (78) attains
@@
$\|S_{tn}-\mathbf{S}_{tn}\|_{1}\leq\frac{1}{n}\sum_{l\leq tn}\|d_{\beta}(S_{l})-d_{\beta}(\mathbf{S}_{l})\|_{1}+\Big{\|}\sum_{l<tn}\Big{(}(S_{l+1}-S_{l})-d_{\beta}(S_{l})\Big{)}\Big{\|}_{1}\,.$ (80)
Since $d_{\beta}$ has bounded Lipschitz coefficient (by some $C_{\beta,q}$), the first sum is at most $\frac{C_{\beta,q}}{n}\sum_{l\leq tn}\|S_{l}-\mathbf{S}_{l}\|_{1}$. By (55), the second sum is a sum of martingale increments (up to an error of $n\cdot O(n^{-2})=O(n^{-1})$), each of which take values in $\{-\frac{1}{n},0,\frac{1}{n}\}$, so by standard martingale concentration estimates (Doob’s maximal inequality and Azuma–Hoeffding bound), with probability $1-o(1)$, the maximum over all $t\leq Tn$ of the second term above is $O(n^{-1/2}\log n)$. Applying the discrete Gronwall inequality, we get

### v2

@@ 4.1 Preliminaries for the Potts Glauber dynamics
    $\exists\delta\geq 0$: $\mathbb{E}_{x_{0}}[X_{t+1}-X_{t}\mid\mathcal{F}_{t}]\leq-\delta$ on $\{X_{t}\geq 0\}$ for all $t\geq 0$.
@@ 4.2 Mixing away from the saddle when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$
We first establish that as soon as the proportion vector has reached an $\omega(n^{-1/2})$ distance from the unstable fixed point of $(m_{*},\frac{1-m_{*}}{q-1},...,\frac{1-m_{*}}{q-1})$ in the first coordinate, with probability going to $1$ as $n\to\infty$, the Glauber dynamics rapidly quasi-equilibrates to the corresponding phase. We need to handle the regimes $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$ and $\beta>\beta_{\mathsf{s}}$ separately, with the latter introducing additional complications; this subsection is focused on the former. Let $\zeta_{n}=o(1)$ be a sequence going to zero sufficiently slowly (chosen after everything else), and define the $q+1$ stable Potts phases when $\beta\in(\beta_{\mathsf{u}},\beta_{\mathsf{s}})$ as
$\Omega^{\mathsf{dis}}$ $=\{\sigma:\|S(\sigma)-(\tfrac{1}{q},...,\tfrac{1}{q})\|_{1}\leq\zeta_{n}\}\,$,
$\Omega^{{\mathsf{ord}},i}$ $=\{\sigma:\|S(\sigma)-(\tfrac{1-m_{\mathsf{r}}}{q-1},...,\tfrac{1-m_{\mathsf{r}}}{q-1},m_{\mathsf{r}},\tfrac{1-m_{\mathsf{r}}}{q-1},...,\tfrac{1-m_{\mathsf{r}}}{q-1})\|_{1}\leq\zeta_{n}\}\,,\qquad\text{and}$
@@
The following allows us to treat the q1q-1 non-dominant coordinates as constant $\frac{1-S_{t,1}}{q-1}$ and approximate the analysis by a 1-dimensional process for the first coordinate.
@@
$\mathbb{P}(\tau_{i,j}^{+}\leq T\wedge\tau_{\varepsilon})\leq 2\exp\Big{(}-\frac{(\frac{A_{2}\log n}{\sqrt{n}}-2/n)^{2}}{8T\cdot(2/n)^{2}}\Big{)}$ $\leq 2\exp\Bigg{(}-\frac{\big{(}\frac{A_{2}\log n}{2\sqrt{n}}\big{)}^{2}}{\frac{2A_{1}n\log n}{n^{2}}}\Bigg{)}$
$=2\exp\Bigg{(}-\frac{A_{2}^{2}\log n}{8A_{1}}\Bigg{)}\,.$
@@
$\left(\frac{d}{dx_{i}}g_{\beta,i}\Big{|}_{x=\hat{S}_{t}}-\frac{d}{dx_{i}}g_{\beta,j}\Big{|}_{x=\hat{S}_{t}}\right)$ $<\frac{qe^{\beta\cdot\hat{S}_{t,i}}}{\sum_{k=1}^{q}e^{\beta\cdot\hat{S}_{t,k}}}$
@@
If $W_{t}\geq 0$, then the first term is at most $\frac{\eta Kn^{-1/2}\log n}{n}$ and the second term is lower order, yielding (60). ∎
@@
$=\frac{1}{n}\cdot[D_{\beta}(m_{*})+(S_{t+\tau_{i},1}-m_{*})\cdot c_{*}+O(|S_{t+\tau_{i},1}-m_{*}|^{2})]+O(n^{-2})$
@@
where the steps in the last line of (67) hold since $s(1+\frac{c_{*}}{16})^{2}<S_{t+\tau_{i},1}-m_{*}<0$, and $s$ is sufficiently small in terms of $c_{*}$ that the second order term is at most $c_{*}(S_{t+\tau_{i},1}-m_{*})/2$, say.
@@
$\big{\|}S(\sigma_{0})-\big{(}m_{\mathsf{r}},\tfrac{1-m_{\mathsf{r}}}{q-1},...,\tfrac{1-m_{\mathsf{r}}}{q-1}\Big{)}\big{\|}_{1}\leq\frac{r}{\sqrt{n}}\quad\text{and}\quad\big{\|}S(\sigma_{0}^{\prime})-\big{(}m_{\mathsf{r}},\tfrac{1-m_{\mathsf{r}}}{q-1},...,\tfrac{1-m_{\mathsf{r}}}{q-1}\Big{)}\big{\|}_{1}\leq\frac{r}{\sqrt{n}}\,.$
@@
$\mathbb{P}\left(\tau_{i+1}\leq\tau_{i}+n,\tilde{\tau}_{i+1}>\min\{\tau_{k+1},\tau_{i}+n\},\tau^{+}_{i+1}>\min\{\tau_{k+1},\tau_{i}+n\}\mid S_{\tau_{i}}\right)$
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\geq 1-3r_{i}-O(n^{-2})\,.$ (70)
@@ 4.3 Getting away from the saddle point when $\beta=\beta_{c}$
$\mathbb{E}[Z_{t+1}-Z_{t}\mid\mathcal{F}_{t}]$ $=n\mathbb{E}[S_{t,1}-S_{t+1,1}\mid\mathcal{F}_{t}]$
$=-d_{\beta}(S_{t})+O(n^{-1})\geq-D_{\beta}(S_{t,1})+O(n^{-1})$
$\geq-C_{1}\gamma n^{-1/2},$
@@
Consider the time and space rescaled process
$\bar{S}_{t}=\sqrt{n}(S_{tn}-(m_{*},\frac{1-m_{*}}{q-1},...,\frac{1-m_{*}}{q-1})).$
By standard results regarding limits of discrete stochastic dynamics as stochastic differential equations (see, e.g., [39]), $\bar{S}_{t}$ converges to the solution of a stochastic differential equation (SDE) $Z_{t}$ on $\mathcal{S}$ with drift and volatility functions with bounded Lipschitz constants. On a compact space, like $\mathcal{S}$, such convergence results only require $O(1)$-Lipschitz bounds on the drift function (which hold in our setting for $g_{\beta}-s$) and moment estimates on the step-wise increments (for which ours are uniformly bounded by $\pm\frac{1}{\sqrt{n}}$).
@@
[…]
@@
for all $\hat{c}_{*}$. Furthermore, $f^{Z}$ is easily checked to be $n$-independent, continuous, monotone, and going to 0 as $\hat{c}_{*}\to\gamma$ and $1$ as $\hat{c}_{*}\to-\gamma$. Therefore, there is a unique $\hat{c}_{*}$ where $f^{Z}$ is the desired $\xi$. Finally, to see that the choice of $\hat{c}_{*}$ is $\gamma$-independent up to $o_{\gamma}(1)$, note that if $\tau_{\gamma}^{-}<\tau_{\gamma}^{+}$, then with probability $1-o_{\gamma}(1)$, also $\tau_{2\gamma}^{-}<\tau_{2\gamma}^{+}$ (by the argument in Lemma 4.6). ∎
@@ 4.4.1 Fixed point analysis of the $q$-dimensional drift function
Now, by (59), $D_{\beta}^{\prime}(x)=0$ has at most two zeros. Aside from $x^{+}$, let $x^{++}$ denote the other zero (if it exists). If $x^{++}$ does not exist or $x^{++}\leq m_{\mathsf{r}}$, then $D_{\beta}^{\prime}(x)<0$ for all $x>m_{\mathsf{r}}$ so $m_{\mathsf{r}}$ is the unique zero in $(1/q,1)$. Now assume $x^{++}>m_{\mathsf{r}}$. By continuity of $D_{\beta}^{\prime}$, $D_{\beta}^{\prime}(x)<0$ for all $x\in(m_{\mathsf{r}},x^{++})$, so by integration $D_{\beta}(x)<0$ for $x\in(m_{\mathsf{r}},x^{++}]$. Moreover, note that $D_{\beta}$ is a monotone function in $(x^{++},1]$ since $D_{\beta}^{\prime}$ does not longer change sign in this interval. Hence, $D_{\beta}$ has no zero on $x^{++},1]$ and $m_{\mathsf{r}}$ is the unique root in $(1/q,1]$. ∎
@@
We first reason that all solutions are of the form of permutations of
$(a,...,a,\frac{1-ka}{q-k},...,\frac{1-ka}{q-k}).$
In order to see this, we suppose by way of contradiction there exists a solution $\bar{s}$ having three distinct values $a,b,c$ appearing in its proportions vector (wlog as $\bar{s}_{1}=a,\bar{s}_{2}=b$ and $\bar{s}_{3}=c$; let $\bar{Z}=\sum_{j}e^{\beta\bar{s}_{j}}$). Then, $e^{\beta a}=a\bar{Z}$ and similarly for $b$ and $c$, and $a\neq b\neq c$. But this is impossible because for any $C,\beta>0$, the equation $e^{\beta x}=Cx$ has at most two solutions on $[0,1]$. When $k=1$, the only possible such solution is the one which has $D_{\beta}=0$ (because the other coordinates are equal, which is where $D_{\beta}=d_{\beta}$), and we can apply Lemma 4.17 to read off that in that case $a=m_{\mathsf{r}}$.
@@
as that would say that such a fixed point is unstable with a drift towards the first coordinate increasing if we perturb in the (1,1,0,,0)(1,-1,0,...,0) direction. To see this, differentiating $g_{\beta,1}$ and plugging in $g_{\beta,1}(\bar{s})=\bar{s}$ and $\bar{s}_{1}=\bar{s}_{2}=\bar{a}$,
@@ 4.4.2 Getting to the stable fixed point once one coordinate dominates
Suppose $S_{0}$ is such that $S_{0,1}\geq S_{0,i}+\delta$ for all $i\geq 2$ for some $\delta>0$. Then, for every $\rho>0$, there exists a time $T=O(n)$ such that $\|S_{T}-(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})\|_{1}\leq\rho$ with probability 1o(1)1-o(1).
@@
therefore it has no closed orbits, and by the bounded Lipschitz constant of $d_{\beta}$, converges exponentially fast to the fixed point in whose basin it is initialized. Moreover, by Lemma 4.18, if initialized with $\mathbf{S}_{0,1}\geq\mathbf{S}_{0,i}+\delta$ for all i=2,,qi=2,...,q for some $\delta>0$, then it is in the basin of attraction of the stable fixed point $(m_{\mathsf{r}},\frac{1-m_{\mathsf{r}}}{q-1},...,\frac{1-m_{\mathsf{r}}}{q-1})$. In particular, for every $\delta>0$, and $\rho>0$, there is a $T=O(1)$ such that under the assumptions on $\mathbf{S}_{0}$ of the lemma, the system of (77) attains
@@
$\|S_{tn}-\mathbf{S}_{tn}\|_{1}\leq\frac{1}{n}\sum_{l<tn}\|d_{\beta}(S_{l})-d_{\beta}(\mathbf{S}_{l})\|_{1}+\Big{\|}\sum_{l<tn}\Big{(}(S_{l+1}-S_{l})-\frac{1}{n}d_{\beta}(S_{l})\Big{)}\Big{\|}_{1}\,.$ (79)
Since $d_{\beta}$ has bounded Lipschitz coefficient (by some $C_{\beta,q}$), the first sum is at most $\frac{C_{\beta,q}}{n}\sum_{l<tn}\|S_{l}-\mathbf{S}_{l}\|_{1}$. By (55), the second sum is a sum of martingale increments (up to an error of $n\cdot O(n^{-2})=O(n^{-1})$), each of which take values in $\{-\frac{1}{n},0,\frac{1}{n}\}$, so by standard martingale concentration estimates (Doob’s maximal inequality and Azuma–Hoeffding bound), with probability 1o(1)1-o(1), the maximum over all $t\leq Tn$ of the second term above is $O(n^{-1/2}\log n)$. Applying the discrete Gronwall inequality, we get